# references.bib

### v2

@@ 3.2. Algorithmic Indirect Discrimination
Where a facially neutral provision, criterion, or practice (PCP) causes disproportionately disadvantages individuals with a protected attribute, it is unlawful indirect discrimination (EqualityAct, 19(1)). Such PCP is discriminatory if it applies to persons with the protected characteristic and puts, or would put, persons with that characteristic at a particular disadvantage when compared to those without such attributes (EqualityAct, 19(2)). There is no requirement that the PCP puts every member of the protected group at a disadvantage, nor the need to prove the reason for the disadvantage (Essop). Further, individuals without the relevant protected characteristic who suffer from the same disadvantage of a discriminatory PCP as those with the protected characteristic can also bring a claim for indirect discrimination (EqualityAct, 19A). The nature of indirect discrimination means there are many potential avenues for proving indirect discrimination in an algorithmic context.
@@ 4.1. Legitimacy of True Differences
Direct discrimination requires formal equality of treatment. It specifically prohibits less favourable treatment of a person based on a protected characteristic, even if there is a true difference in their risk, ability, or merit. In addition to prohibiting prejudice, direct discrimination acknowledges that there are certain protected characteristics that should not be used as the basis for less favourable treatment even though that characteristic may be relevant. Although gender, pregnancy, certain disabilities, among other things may be relevant to a person’s job performance, the law prohibits discrimination on those grounds (khaitan2015). For example, until 2012 insurers could use sex as a determining factor in risk assessments based on actuarial and statistical data. However, the European Court of Justice overturned this rule (TestAchats), arguing that the use of sex as a risk factor should not result in differences in individuals’ insurance premiums and benefits (EA2012). Instead of using sex as a variable, insurers need to consider information that is not protected and closer to predicting the true question of risk. Direct discrimination “allows only carefully defined distinctions and otherwise expects symmetry” (EandJFS). These defined distinctions are important. For instance, insurance decisions that might otherwise be construed as discriminatory – specifically concerning gender reassignment, marriage, civil partnership, pregnancy, and sex discrimination – are permissible if they are based on reliable actuarial data and executed reasonably (EqualityAct, Sch 9 s 20). Similarly, financial services can “use age as a criterion for pricing risk, as it is a key risk factor associated with for example, medical conditions, ability to drive, likelihood of making an insurance claim and the ability to repay a loan” (agediscrimnotes, para 7.6). Similar statutory exemptions are found in anti-discrimination laws in the EU (EU2002, art 2), Australia (AUSDA, s 30-47), Canada (CanadaHRA, s 15), New Zealand (NZHRA, s 24-60) and South Africa (SADL, s 14). These exemptions legally recognise that certain group distinctions, particularly those involving risk assessment, are relevant and necessary for the equitable operation of such services.
@@
On the other hand, indirect discrimination recognises that formal neutrality or even the application of formal equality can create disadvantages to protected groups. Avoiding indirect discrimination may sometimes require providing opportunities in a manner that only discriminates when accounting for legitimate individual and group differences – aiming for substantive equality (vos2020; xenidis2020; hellman20202; xiang2021; wachter20211). Consider a scenario where a store manager imposes a minimum height requirement for all employees, justifying the policy by the need for everyone to equally be able to access supplies stored on high shelves. Although the rule may appear neutral, it would, in practice, disproportionately disadvantage women, whose average heights are lower than those of men. Indirect discrimination recognises the reality of the true height differences between certain groups.
@@ 4.2. Ground Truth, True Data Generating Process and Estimation
Similar to related work on measurement bias and construct validity (jacobs2021), we explore a fundamental and overlooked limitation of legal frameworks in algorithmic discrimination: law is predicated on an individualistic model that often requires a threshold of identifiable harm before legal action can be triggered. It is not inherently designed to work with approximations or mitigating systemic level harms. We outline a critical intervention that emerges from comparing true underlying risks and estimation of that risk. Therefore, an important aspect from the legal perspective that is overlooked in the algorithmic fairness literature, but is a standard theoretical framework in statistics, the distinction between a true DGP and the estimated model $\hat{p}(y|x)$. To formalise, we assume that there exists a true DGP, $D\sim p(y,x)$, where $D_{i}=(y_{i},x_{i})$. Further, we use $p(y|x^{\text{true}}_{i})$ to denote the true probability (pmf) for individual $i$, given the true features $x^{\text{true}}_{i}$.
@@
We make multiple observations on the role of the true model and its use in connecting predictive modelling and legal reasoning. First, understanding the limits of predictive models is crucial to explore inherent uncertainties and limitations in predictions. The true model is, in practice, never observed or known. When developing $\hat{p}(y|x)$, the target is often to select the model with the best predictive performance, which is closely connected to the true DGP (shao1993linear; bernardo1994bayesian; vehtari2002bayesian; vehtari2012survey; vehtari2017practical). The true model may include features in $x^{\text{true}}_{i}$ that are not observed in the data, sometimes referred to as an $\mathcal{M}$-open setting, i.e., when the true model is not included in the set of candidate models (bernardo1994bayesian; vehtari2012survey).
@@
Second, we assume that $p(y|x_{i})$ is a probability distribution over $\mathcal{Y}$, introducing some level of aleatoric uncertainty in the true underlying process (o_hagan2004; Hullermeier2021; tahir2023). This means that perfect prediction of $y_{i}$ may not be possible, even with knowledge of the true DGP. The distinction between aleatoric and epistemic uncertainty is important from a legal perspective. The reason is simple: the uncertainty coming from estimation and modelling is the (legal) responsibility of the modeller, while the aleatoric uncertainty can instead be considered a true underlying general risk.
@@
Third, the true DGP connects to judicial legal reasoning. Courts must engage theoretically with legal and normative conceptions of what constitutes unlawful discrimination and what is justifiable. Judges consider legitimacy, proportionality, and necessity when evaluating actions, and hypothetical alternatives, that led to discriminatory treatment or outcomes. Although case law does not always represent ground truth, courts can operate as an oracle. It may not pinpoint what the perfect decision should have been, but courts will engage in a similar theoretical process of reasoning about the decision-making process to the true DGP to understand whether the actions were justified or unlawful. We explain legal reasoning within this framework throughout the paper and in a real-world case on unlawful discrimination in algorithmic decision-making (Section 7).
@@ 4.3. Estimation Parity
Algorithmic fairness literature often assumes the absence of estimation error (see e.g., hardt2016) or assumes that the true causal structure is known (Zhang2017; kilbertus2018; chiappa2019; carey2022causal). In practice, this is rarely the case. Hence, it is crucial, both practically and legally, to distinguish between the true underlying probabilities $\boldsymbol{\pi}_{i}$ and the estimated probabilities $\hat{\boldsymbol{\pi}}_{i}$. While disparities in the true underlying probability may sometimes be legitimate or justified (Section 4.2), introducing an estimation error that disadvantages individuals based on protected attributes invokes discrimination liability.
As the model will try to approximate the true DGP, modellers’ expectations are difficult to ascertain. Law is unlikely to set a deterministic standard that any adverse effects of estimation will make a modeller liable, yet the modeller should try to approximate the true model as well as possible (see akaike1973information; white1982maximum; vehtari2012survey; vehtari2017practical, for discussions on model misspecification). Where an estimation disparity reaches a threshold for discriminatory effects, the legal evaluation would require analysing the steps taken to test and mitigate estimation disparity (even though the intent is immaterial).
The potential bias in training data presents a risk that the estimation model will introduce bias against individuals with protected attributes (Section 6.2). Historical discriminatory lending practices, for example, could be perpetuated through biased training data (black1978; rice2013; sargeant2023). Such biased estimations may introduce biased outcomes that are not reflective of true differences, potentially leading to discriminatory outcomes. Therefore, we introduce “Conditional Estimation Parity” to formalise the legal context of estimation.
@@
Reducing the error in Eq. 4 is expected to diminish the risk of conditional estimation disparity. However, assessing conditional estimation parity is complex due to inherent challenges in evaluating estimation error. It is crucial to examine both mathematical and legal causal theories of why certain differences are legitimate bases to make classification distinctions (kiviat2023). Assuming a model action appears to be prohibited by the rule articulation function, the legal analysis shifts to whether liability can be proven.
@@ 5.1. Statistical Disparities and Prima Facie Discrimination
To initiate a claim for direct or indirect discrimination, a claimant must establish a prima facie case (EWCA2005; EqualityAct, s 136).
In direct discrimination, the claimant must show the alleged discrimination explicitly referred to a protected characteristic or exact proxy; the protected characteristic has to be the reason for the treatment (court2018; court2017). For indirect discrimination, sufficient evidence must be produced to identify the PCP, identify the protected group, and show that the PCP places the protected group at a particular disadvantage when compared to those without such attribute (EWCA2005; Essop). Although courts do not always defer to statistical evidence (court2012), it is commonly used and will likely be essential in an algorithmic context.
Some algorithmic fairness papers have written that the US law defines explicit statistical thresholds to define a prima facie case, such as the four-fifths rule in employment discrimination law (romei2014; feldman2015; zafar2017). However, this conflates the legal position of these tests, which are certainly not rigid definitions of differential outcomes (EEOC; Watkins2024; Harned2020). In any event, UK law is more flexible than the US on thresholds for statistical significance which are often resisted by courts to avoid excessive dependence on data (Seymour; sanchez2020). Statistical significance depends on the context of the comparison (Essop; Seymour), and smaller disparities are less likely to trigger legal inquiry under anti-discrimination laws (Essop).
@@
(6) $\omega=||\mathbb{E}_{x}\left[\epsilon_{i}\mid x_{l},x_{p}\right]-\mathbb{E}_{x}\left[\epsilon_{i}\mid x_{l}\right]\,||_{2},$
where $||\cdot||_{2}$ is the euclidean norm. This target generalises the idea of minimising conditional statistical parity. If we assume true conditional statistical parity, i.e.
(7) $\mathbb{E}_{x}\left[p(y_{i}|x_{i}^{\text{true}})\mid x_{l},x_{p}\right]=\mathbb{E}_{x}\left[p(y_{i}|x_{i}^{\text{true}})\mid x_{l}\right]\,,$
@@
Hence, if true statistical parity does not hold, it is explained by true differences between groups. If there is a true difference, such as age in financial services, forcing conditional statistical parity could harm the protected group, most likely resulting in unlawful discrimination. This result aligns with previous observations about the risks of forcing parity metrics (corbett2017; xiang2021; hertweck2021). Courts may need to be more flexible in the type of statistical data they consider to establish a prima facie case by considering non-comparative adverse effects in their assessment. Therefore, deferring to conditional estimation parity provides an avenue for a contextually informed assessment.
@@ 5.2. Legal Causation for Unlawful Discrimination
To lawyers, causation is the relationship between an act and its effect, which requires two questions: (1) factually, but for the act, would the consequences have occurred; (2) is the act a substantial cause of the consequence to apply legal responsibility. We are concerned with the first question. Direct discrimination “requires a causal link between the less favourable treatment and the protected characteristic”; indirect discrimination “requires a causal link between the PCP and the particular disadvantage suffered by the group and individual [sharing a particular protected characteristic]” (Essop).
@@
There is ongoing debate about whether legal causation poses an insurmountable obstacle in proving algorithmic discrimination, given the technical difficulties in establishing direct causal links between a protected characteristic and the resulting outcome. Some scholars highlight distinct and higher standard of causation under US law (hurley2017; barocas2016) , while some focus narrowly on statistical causality and correlations, which do not always align with legal causation (xenidis2020; williams2018). While it is true even in the UK that “a correlation is not the same as a causal link” (Essop), courts have shown a functional flexibility in the application of causation in order to navigate complex human decision-making processes to determine legal causation (EandJFS), which can be even more intricate or opaque than algorithmic processes.
@@
In an algorithmic context, this causal link requires asking whether $i$ would have received the same action or decision $a$, but for their protected attribute $x_{p}$ or the PCP that indirectly relates to their protected attribute $x_{p}$ (lords1989; lords1990). From a decision-theoretic perspective, the protected attribute $x_{p}$ can affect the decision $a$ either through the utility function $u(a,y)$ or through the model $\hat{p}(y|x)$. Discrimination may occur if the utility function in Eq. 1 differs for different groups defined by the protected attribute. This concept parallels taste-based discrimination, where a decision-maker’s subjective preferences or prejudices against a group lead to differences in outcomes (becker1957). In this context, the utility function $u(a,y)$ unjustifiably disfavours a group based on protected attributes $x_{p}$. Such a difference would mean that an individual or whole group with a protected attribute is treated less favourably than those without a protected attribute given the same model $\hat{p}(y|x)$. Such a difference in the utility function would risk unlawful discrimination. Specifically, if $u(a,y)$ is changed for different persons, either directly based on a protected attribute or indirectly has the effect of disproportionately disadvantaging a group with a protected characteristic without justification.
@@
Differing $\hat{p}(y|x)$, on the other hand, would mean that there is a legal causation between the decision $a$ and $x_{p}$. The differing $\hat{p}(x|y)$ is analogous to statistical discrimination, where group-level statistics are used as proxies for individual characteristics due to imperfect information (arrow1971; phelps1972). This might either be motivated by true differences or a result of conditional estimation disparity. In the latter case, this might be a case of legal causation, i.e., that the model is poor, and hence, the modelling has resulted in disadvantaging a protected group. Therefore, we can view the legal causal structure of $\hat{p}(y|x)$ as central to avoiding unlawful discrimination. However, not considering legal causation could lead to conditional estimation disparity, and potentially result in unlawful discrimination.
@@
Legal causation focuses on the causal link between $x_{p}$ and the decision $a$. Additionally, legal causation is less formal than common definitions of causal effects in ML. Courts, at least outside of the US, are effects-orientated, and a wide range of forms of a “legal causal link” could be identified (shin2013; khaitan2015). Much of the causal-based fairness literature formulates “causation” on the true causal model structure in $\hat{p}(y|x)$, i.e., the study of the causal effect of $x$, due to outside interventions on $y$ (Pearl2010; barocas2016; zhang2018fairness; carey2022causal). However, this formulation is not the same as that of legal causation.

### v3

@@ 3.2. Algorithmic Indirect Discrimination
Where a facially neutral provision, criterion, or practice (PCP) disproportionately disadvantages individuals with a protected attribute, it is unlawful indirect discrimination (United Kingdom Parliament, 2010, s 19(1)). The UK Supreme Court (UKSC) explains that it “aims to achieve equality of results in the absence of such justification” (United Kingdom Supreme Court, 2017b, para 25). Such PCP is discriminatory if it applies to persons with the protected characteristic and puts, or would put, persons with that characteristic at a particular disadvantage when compared to those without such attributes (United Kingdom Parliament, 2010, s 19(2)). There is no requirement that the PCP puts every member of the protected group at a disadvantage, nor the need to prove the reason for the disadvantage (United Kingdom Supreme Court, 2017b). Further, even individuals without the relevant protected characteristic who suffer from the same disadvantage of a discriminatory PCP as those with the protected characteristic can also bring a claim for indirect discrimination (also known as “discrimination by association”) (established in EU case law European Court of Justice, 2015),(adopted in the UK in 2024, United Kingdom Parliament, 2010, s 19A). Therefore, the definition of indirect discrimination is broader under UK law than in the US and, as will be discussed below, the thresholds for proving liability is lower in the UK than under US doctrine. The scope of indirect discrimination means there are many potential avenues for it to arise in an algorithmic context.
@@ 4.1. Legitimacy of True Differences
Direct discrimination requires formal equality of treatment. It specifically prohibits less favourable treatment of a person based on a protected characteristic, even if there is a true difference in their risk, ability, or merit. In addition to prohibiting prejudice, direct discrimination acknowledges that there are certain protected characteristics that should not be used as the basis for less favourable treatment even though that characteristic may be relevant. Although gender, pregnancy, certain disabilities, among other things may be relevant to a person’s job performance, the law prohibits discrimination on those grounds (Khaitan, 2015). For example, until 2012 insurers could use sex as a determining factor in risk assessments based on actuarial and statistical data. However, the European Court of Justice overturned this rule (European Court of Justice, 2011), arguing that the use of sex as a risk factor should not result in differences in individuals’ insurance premiums and benefits (United Kingdom Parliament, 2012b). Instead of using sex as a variable, insurers need to consider information that is not protected and closer to predicting the true question of risk. Direct discrimination “allows only carefully defined distinctions and otherwise expects symmetry” (United Kingdom Supreme Court, 2009). These defined distinctions are important. For instance, insurance decisions that might otherwise be construed as discriminatory – specifically concerning gender reassignment, marriage, civil partnership, pregnancy, and sex discrimination – are permissible if they are based on reliable actuarial data and executed reasonably (United Kingdom Parliament, 2010, Sch 9 s 20). Similarly, financial services can “use age as a criterion for pricing risk, as it is a key risk factor associated with for example, medical conditions, ability to drive, likelihood of making an insurance claim and the ability to repay a loan” (United Kingdom Parliament, 2012a, para 7.6). Similar statutory exemptions are found in anti-discrimination laws in the EU (European Parliament, 2002, art 2), Australia (Australian Parliament, 1984, s 30-47), Canada (Canadian Parliament, 1985, s 15), New Zealand (New Zealand Parliament, 1993, s 24-60) and South Africa (South African Parliament, 2000, s 14). These exemptions legally recognise that certain group distinctions, particularly those involving risk assessment, are relevant and necessary for the equitable operation of such services.
@@
On the other hand, indirect discrimination recognises that formal neutrality or even the application of formal equality can create disadvantages to protected groups. Avoiding indirect discrimination may sometimes require providing opportunities in a manner that only discriminates when accounting for legitimate individual and group differences – aiming for substantive equality (Vos, 2020; Xenidis, 2020; Wachter et al., 2021b). Consider a scenario where a store manager imposes a minimum height requirement for all employees, justifying the policy by the need for all employees to equally be able to access supplies stored on high shelves. Although the rule may appear neutral, it would, in practice, disproportionately disadvantage women, whose average heights are lower than those of men. Indirect discrimination recognises the true height differences between certain groups.
@@ 4.2. Ground Truth, True Data Generating Process and Estimation
Similar to related work on measurement bias and construct validity (Jacobs and Wallach, 2021), we explore a fundamental and overlooked limitation of legal frameworks in algorithmic discrimination: law is predicated on an individualistic model that often requires a threshold of identifiable harm before legal action can be triggered. It is not inherently designed to work with approximations or mitigating systemic level harms. We outline a critical intervention that emerges from comparing true underlying risks and estimation of that risk. Therefore, an important aspect from the legal perspective that is overlooked in the algorithmic fairness literature, but is a standard theoretical framework in statistics, the distinction between a true DGP and the estimated model $\hat{p}(y|x)$. To formalise, we assume that there exists a true DGP, $D\sim p(y,x)$, where $D_{i}=(y_{i},x_{i})$. Further, we use $p(y|x^{\text{true}}_{i})$ to denote the true probability (pmf) for individual $i$, given the true features $x^{\text{true}}_{i}$.
We make multiple observations on the role of the true model and its use in connecting predictive modelling and legal reasoning.
@@
First, understanding the limits of predictive models is crucial to explore inherent uncertainties and limitations in predictions. The true model is, in practice, never observed or known. When developing $\hat{p}(y|x)$, the target is often to select the model with the best predictive performance, which is closely connected to the true DGP (Shao, 1993; Bernardo and Smith, 1994; Vehtari and Lampinen, 2002; Vehtari and Ojanen, 2012; Vehtari et al., 2017). The true model may include features in $x^{\text{true}}_{i}$ that are not observed in the data, sometimes referred to as an $\mathcal{M}$-open setting, i.e., when the true model is not included in the set of candidate models (Bernardo and Smith, 1994; Vehtari and Ojanen, 2012).
@@
Second, we assume that $p(y|x_{i})$ is a probability distribution over $\mathcal{Y}$, introducing some level of aleatoric uncertainty in the true underlying process (O’Hagan, 2004; Hüllermeier and Waegeman, 2021; Tahir et al., 2023). This means that perfect prediction of $y_{i}$ may not be possible, even with knowledge of the true DGP. The distinction between aleatoric and epistemic uncertainty is important from a legal perspective. The reason is simple: the uncertainty coming from estimation and modelling is the (legal) responsibility of the modeller, while the aleatoric uncertainty can instead be considered a true underlying general risk.
@@
Third, the true DGP connects to judicial legal reasoning. Courts must engage theoretically with legal and normative conceptions of what constitutes unlawful discrimination and what is justifiable. Judges consider legitimacy, proportionality, and necessity when evaluating actions, and hypothetical alternatives, that led to discriminatory treatment or outcomes. Although case law does not always represent ground truth, courts can operate as an oracle. It may not pinpoint what the perfect decision should have been, but courts will engage in a similar theoretical process of reasoning about the decision-making process to the true DGP to understand whether the actions were justified or unlawful. We explain legal reasoning within this framework throughout the paper and in a case on unlawful discrimination in algorithmic decision-making (Section 7).
@@ 4.3. Estimation Parity
Algorithmic fairness literature often assumes the absence of estimation error (see e.g., Hardt et al., 2016) or assumes that the true causal structure is known (Zhang et al., 2017; Kilbertus et al., 2018; Chiappa, 2019; Carey and Wu, 2022). In practice, this is rarely the case. Hence, it is crucial, both practically and legally, to distinguish between the true underlying probabilities $\boldsymbol{\pi}_{i}$ and the estimated probabilities $\hat{\boldsymbol{\pi}}_{i}$. While disparities in the true underlying probability may sometimes be legitimate or justified (Section 4.2), introducing an estimation error that disadvantages individuals based on protected attributes invokes discrimination liability.
As the model will try to approximate the true DGP, modellers’ expectations are difficult to ascertain. Law is unlikely to set a deterministic standard that any adverse effects of estimation will make a modeller liable, yet the modeller should try to approximate the true model as well as possible (see Akaike, 1973; White, 1982; Vehtari and Ojanen, 2012; Vehtari et al., 2017, for discussions on model misspecification). Where an estimation disparity reaches a threshold for discriminatory effects, the legal evaluation would require analysing the steps taken to test and mitigate estimation disparity (even though the intent is immaterial).
The potential bias in training data presents a risk that the estimation model will introduce bias against individuals with protected attributes (Section 6.2). Historical discriminatory lending practices, for example, could be perpetuated through biased training data (Black et al., 1978; Rice and Swesnik, 2013; Sargeant, 2023, 2025). Such biased estimations may introduce biased outcomes that are not reflective of true differences, potentially leading to discriminatory outcomes. Therefore, we introduce “Conditional Estimation Parity” to formalise the legal context of estimation.
@@
Reducing the error in Eq. 4 is expected to diminish the risk of conditional estimation disparity. However, assessing conditional estimation parity is complex due to inherent challenges in evaluating estimation error. It is crucial to examine both mathematical and legal causal theories of why certain differences are legitimate bases to make classification distinctions (Kiviat, 2023). Assuming a model action appears to be prohibited by the rule articulation function, the legal analysis shifts to whether liability can be proven. We now examine the basis for identifying statistical disparities and legitimate differentiation in the context of unlawful discrimination.
@@ 5.1. Statistical Disparities and Prima Facie Discrimination
To initiate a claim for direct or indirect discrimination, a claimant must establish a prima facie case (England and Wales Court of Appeal, 2005; United Kingdom Parliament, 2010, s 136).
In direct discrimination, the claimant must show the alleged discrimination explicitly referred to a protected characteristic or exact proxy; the protected characteristic has to be the reason for the less favourable treatment (United Kingdom Supreme Court, 2018, 2017c). For indirect discrimination, sufficient evidence must be produced to identify the PCP, identify the protected group, and show that the PCP places the protected group at a particular disadvantage when compared to those without such attribute (England and Wales Court of Appeal, 2005; United Kingdom Supreme Court, 2017b). Although courts do not always defer to statistical evidence (United Kingdom Supreme Court, 2012), it is commonly used and will likely be essential in an algorithmic context.
Some algorithmic fairness papers have written that the US law defines explicit statistical thresholds to define a prima facie case, such as the four-fifths rule in employment discrimination law (Romei and Ruggieri, 2014; Feldman et al., 2015; Zafar et al., 2017). However, this conflates the legal position of these tests, which are certainly not rigid definitions of differential outcomes (United States Government, 2023; Watkins and Chen, 2024; Harned and Wallach, 2020). In any event, UK law is more flexible than the US on thresholds for statistical significance which are often resisted by courts to avoid excessive dependence on data (United Kingdom House of Lords, 2000; Sánchez-Monedero et al., 2020). Statistical significance depends on the context of the comparison (United Kingdom Supreme Court, 2017b; United Kingdom House of Lords, 2000), and smaller disparities are less likely to trigger legal inquiry under anti-discrimination laws (United Kingdom Supreme Court, 2017b).
@@
(6) $\omega=|\mathbb{E}_{x}\left[\epsilon\mid x_{l},x_{p}\right]-\mathbb{E}_{x}\left[\epsilon\mid x_{l}\right]\,|,$
where ||$\cdot$|| ⋅ | denotes the absolute value. This target generalises the idea of minimising conditional statistical parity. If we assume true conditional statistical parity, i.e.
(7) $\mathbb{E}_{x}\left[p(y|x^{\text{true}})\mid x_{l},x_{p}\right]=\mathbb{E}_{x}\left[p(y|x^{\text{true}})\mid x_{l}\right]\,,$
@@
Hence, if true statistical parity does not hold, it is explained by true differences between groups. If there is a true difference, such as age in financial services, forcing conditional statistical parity could harm the protected group, most likely resulting in unlawful discrimination. This result aligns with previous observations about the risks of forcing parity metrics (Corbett-Davies et al., 2017; Xiang, 2021; Hertweck et al., 2021; Sargeant, 2025). Courts may need to be more flexible in the type of statistical data they consider to establish a prima facie case by considering non-comparative adverse effects in their assessment. Therefore, deferring to conditional estimation parity provides an avenue for a contextually informed assessment.
@@ 5.2. Legal Causation for Unlawful Discrimination
To lawyers, causation is the relationship between an act and its effect, which requires two questions: (1) factually, but for the act, would the consequences have occurred; (2) is the act a substantial cause of the consequence to apply legal responsibility. We are concerned with the first question. Direct discrimination “requires a causal link between the less favourable treatment and the protected characteristic”; indirect discrimination “requires a causal link between the PCP and the particular disadvantage suffered by the group and individual [sharing the protected characteristic]” (United Kingdom Supreme Court, 2017b).
@@
There is ongoing debate about whether legal causation poses an insurmountable obstacle in proving algorithmic discrimination, given the technical difficulties in establishing direct causal links between a protected characteristic and the resulting outcome. Some scholars highlight distinct and higher standard of causation under US law (Hurley and Adebayo, 2017; Barocas and Selbst, 2016, discussing the heightened “robust causality requirement” introduced in US Supreme Court case, Inclusive Communities), while some focus narrowly on statistical causality and correlations, which do not always align with legal causation (Xenidis, 2020; Williams et al., 2018). While it is true even in the UK that “a correlation is not the same as a causal link” (United Kingdom Supreme Court, 2017b), courts have shown a functional flexibility in the application of causation in order to navigate complex human decision-making processes to determine legal causation (United Kingdom Supreme Court, 2009), which can be even more intricate or opaque than algorithmic processes.
@@
In an algorithmic context, this causal link requires asking whether $i$ would have received the same action or decision $a$, but for their protected attribute $x_{p}$ or the PCP that indirectly relates to their protected attribute $x_{p}$ (United Kingdom House of Lords, 1989, 1990). From a decision-theoretic perspective, the protected attribute $x_{p}$ can affect the decision $a$ either through the utility function $u(a,y)$ or through the model $\hat{p}(y|x)$. Discrimination may occur if the utility function in Eq. 1 differs for different groups defined by the protected attribute. This concept parallels taste-based discrimination, where a decision-maker’s subjective preferences or prejudices against a group lead to differences in outcomes (Becker, 1957). In this context, the utility function $u(a,y)$ unjustifiably disfavours a group based on protected attributes $x_{p}$. Such a difference would mean that an individual or whole group with a protected attribute is treated less favourably than those without a protected attribute given the same model $\hat{p}(y|x)$. Such a difference in the utility function would risk unlawful discrimination. Specifically, if $u(a,y)$ is changed for different persons, either directly based on a protected attribute or it indirectly has the effect of disproportionately disadvantaging a group with a protected characteristic without justification. A detailed legal assessment would consider the specifics of the case to determine legal causation.
@@
Differing $\hat{p}(y|x)$, on the other hand, would mean that there is an indirect causation between the decision $a$ and $x_{p}$. The differing $\hat{p}(x|y)$ is analogous to statistical discrimination, where group-level statistics are used as proxies for individual characteristics due to imperfect information (Arrow, 1971; Phelps, 1972). This might either be motivated by true differences or a result of conditional estimation disparity. In the latter case, this might be a case of legal causation, i.e., that the model is poor, and hence, the modelling has resulted in disadvantaging a protected group. Therefore, we can view the legal causal structure of $\hat{p}(y|x)$ as central to avoiding unlawful discrimination. However, not considering legal causation could lead to conditional estimation disparity, and potentially result in unlawful discrimination.
@@
Legal causation focuses on the causal link between $x_{p}$ and the decision $a$. Additionally, legal causation is less formal than common definitions of causal effects in ML. Courts, at least outside of the US, are effects-orientated, and a wide range of forms of a “legal causal link” could be identified (Shin, 2013; Khaitan, 2015). Much of the causal-based fairness literature formulates “causation” on the true causal model structure in $\hat{p}(y|x)$, i.e., the study of the causal effect of $x$, due to outside interventions on $y$ (Pearl, 2010; Barocas and Selbst, 2016; Zhang and Bareinboim, 2018; Carey and Wu, 2022). However, this formulation is not the same as that of legal causation.